OpenStudy (anonymous):

if tan(2pi/3)sin(3pi/2-x)=1 what is cos(2x)?

OpenStudy (kc_kennylau):

well, try to simplify sin(3pi/2-x) first :)

OpenStudy (kc_kennylau):

and of course find the value of tan(2pi/3) first :)

OpenStudy (anonymous):

Hi!! \(\huge\color{blue}{{Welcome~To~Open~Study}!!!!!!}\)

OpenStudy (kittiwitti1):

This?\[\text{if }\tan{\frac{2\pi}{3}}\sin{\frac{3\pi}{2-x}}=1\text{ what is}\cos{2x}\]

OpenStudy (anonymous):

yes

OpenStudy (kittiwitti1):

Hm, alright then

OpenStudy (kc_kennylau):

@kittiwitti1 I believe that it is \(\displaystyle \sin\left(\frac{3\pi}2-x\right)\) instead

OpenStudy (kittiwitti1):

I think so too @kc_kennylau but that is why I asked :)

OpenStudy (anonymous):

@kc_kennylau you're right sry!

OpenStudy (kittiwitti1):

xD \[\text{Well, tangent is also equal to sine divided by cosine:}\tan{\theta}=\frac{\sin{\theta}}{\cos{\theta}}\]

OpenStudy (kittiwitti1):

...did everyone else leave ._.

OpenStudy (anonymous):

@kittiwitti1 tnx

OpenStudy (kittiwitti1):

Lol... Ok. We have two radian values, 2pi/3 and (3pi/2)-x\[\text{What do }\frac{2\pi}{3}\text{ and }\frac{3\pi}{2}\text{ equal in degrees?}\]

OpenStudy (kittiwitti1):

it means 2π/3=a° find a what did you get? :)

OpenStudy (kittiwitti1):

*2pi/3 = a degrees, find a

OpenStudy (kittiwitti1):

http://prntscr.com/67ip8y Yes you got it right :)

OpenStudy (anonymous):

@kittiwitti1 2pi/3=120

OpenStudy (kittiwitti1):

^-^

OpenStudy (kittiwitti1):

So now can you solve with the degree values to find x? :)

OpenStudy (anonymous):

@kittiwitti1 thanks

OpenStudy (kittiwitti1):

You're welcome~ but are you ok now or do you need more help :p

OpenStudy (anonymous):

@kittiwitti1 i'm ok and i found the answer

OpenStudy (kittiwitti1):

Okay :)

OpenStudy (kittiwitti1):

Don't forget to close the question ^-^